OpenStudy (compassionate):

My friend doesn't understand this equation and %100 sure that she knows the right answer: 48÷2(9+3) = Go, Mathbro's, show her your knowledge.

OpenStudy (vishweshshrimali5):

She should know the rule of VBODMAS

OpenStudy (vishweshshrimali5):

First solve verniculum, then bracket, then off, then division, then multiplication, then addition and at last subtraction

OpenStudy (compassionate):

Give it a shot.

OpenStudy (vishweshshrimali5):

See, first in this question what is first bracket

OpenStudy (lgbasallote):

i think you mean vinculum. the thingy like \(\overline{x+3}\)

OpenStudy (vishweshshrimali5):

solve what is in the bracket

OpenStudy (vishweshshrimali5):

so 9+3 = 12

OpenStudy (vishweshshrimali5):

Yes @lgbasallote

OpenStudy (vishweshshrimali5):

So, the equation becomes 48 / 12

OpenStudy (vishweshshrimali5):

Now what is left is division so divide 48 by 12 u get 4 which is ur answer

OpenStudy (vishweshshrimali5):

Got that @Compassionate

OpenStudy (lgbasallote):

you forgot the 2

OpenStudy (vishweshshrimali5):

Woops sorry

OpenStudy (compassionate):

48÷2(9+3) 48÷2(9+3) 24 (9 + 3) 24 x 12 288

OpenStudy (lgbasallote):

interesting is if you would do it like \[\frac{48}{2(9+3)}\]ir\[\frac{48}{2}(9+3)\]

OpenStudy (vishweshshrimali5):

Yes because division comes first

OpenStudy (vishweshshrimali5):

@lgbasallote when the question is given in the format @Compassionate has given then we use the VBODMAS rule to clear all such doubts

OpenStudy (lgbasallote):

shouldnt it be VBBPMAS

OpenStudy (lgbasallote):

there isnt a mathematical operation "off" is there?

OpenStudy (vishweshshrimali5):

Yes there is

OpenStudy (lgbasallote):

oh exponent

OpenStudy (vishweshshrimali5):

I meant "of"

OpenStudy (vishweshshrimali5):

Nope

OpenStudy (lgbasallote):

well you dont have exponent in your "VBODMAS"

OpenStudy (vishweshshrimali5):

I meant yes for exponents and nope for your comment that there doesn't exist any

OpenStudy (compassionate):

Now it begins.

OpenStudy (vishweshshrimali5):

@lgbasallote of is for exponents

OpenStudy (lgbasallote):

that's an odd way of putting it lol

OpenStudy (kainui):

Which problem are we solving here people? Hello? \[\frac{ 48 }{ 2(9+3) } or \frac{ 48 }{ 2 }(9+3)\]

OpenStudy (kainui):

Also consider which answer is harder to get a whole number for? The one with an answer of 2 or 288? Generally problems are made.

OpenStudy (vishweshshrimali5):

@Kainui what if the same question comes in your exam........ u r not going to write that what do u mean to say................ to remove such problems the rule of VBODMAS is followed

OpenStudy (kainui):

The rule IS followed. What you have to understand is that the left answer is correct. When I wrote it like that it becomes immediately obvious why. See in the left formula you have 2(9+3) whereas in the right formula it is interpreted as .5(9+3) which is clearly wrong. So to clarify, 2 is the correct answer.